An increase of 10K in temperature rarely doubles the kinetic energy of particles
but this much increase in temperature may be enough to double the rate of reaction.2?

Grade:9

2 Answers

Sunil Kumar FP
9 years ago
an increse of 10k is not enough to double the kinetic energy because the molecular spped is not going to change significantly wheres at rate of reaction doubles
consider for example
T1=298k
T2=308K
Ea=52.8kj
from formula=logk2/k1=Ea/2.303R(1/T2-1/T2)
k2/k1=2
thus the rate of the reaction doubles
Priyanshi soni
4 years ago
Because on increasing temperature by 10K the fraction of molecules  having effective collision is increased by double hence, reaction rate get double.